OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x 2 sin x cos2x + cos x 2 sin x cos2x + sin3x sin x cos2x - sin3x + cos3x 2 cos2x sin x + sin x - 2 sin3x

OpenStudy (bahrom7893):

Sin(3x) = Sin(2x+x)

OpenStudy (jhannybean):

sin (3x) = sin (2x +x) sin(2x +x ) = sin (2x)cos(x) + cos(2x) sin(x)

OpenStudy (bahrom7893):

Just use double/triple angle formulas, and you're good to go.

OpenStudy (jhannybean):

sin(2x) cos(x) + cos(2x) sin(x) sin(2x) = 2 sin(x) cos(x) cos(2x) = 1 - 2 sin^(2) (x)

OpenStudy (jhannybean):

So... simplify sin(2x) cos(x) + cos(2x) sin(x) \(\large \left(2sin(x)cos(x)\right)\left(cos(x)\right) + \left(1-2sin^{2}(x)\right)\left(sin(x)\right)\)

OpenStudy (jhannybean):

Now simplify this and you'll get your answer :)

OpenStudy (anonymous):

so the last one?

OpenStudy (jhannybean):

\(\large \left(2sin(x)cos^2(x)\right) + \left(sin(x)-2sin^3(x)\right)\)

OpenStudy (jhannybean):

Yesss :D